OpenStudy (anonymous):

Solve: 6x - 40/ 2 = 4 a. x = 16/3 b. x = 22/3 c. x = 6 d. x = 8

OpenStudy (anonymous):

(6x - 40)/2 = 4 becomes 6x/2 - 40/2 = 4 which is 3x - 20 = 4. This simplifies to: 3x = 24 and x = 8 by basic algebra. If the equation is 6x - 40/2 = 4, then you have 6x - 20 = 4, and 6x = 24 by basic algebra and x = 4. Hope this helps. If it does, please choose good answer. Visit http://www.tutorsean.net for free math videos.